### v1

# A Simple Model of Current Ramp-Up and Ramp-Down in Tokamaks

Abstract
[…]
I Introduction
Richard Fitzpatrick’s 2026 article Fitzpatrick:2026 on current ramp-up and ramp-down in tokamaks, confusingly coupled a number of effects that are in principle separable. A disentanglement, Section II, shows his fundamental physics errors in the definition and evolution of the poloidal flux in tokamaks. In particular, Fitzpatrick:
  1. 1.
    Did not explicitly include the changes in the flux in the central solenoid in his analysis, which is the only flux that is a directly controllable function of time.
  2. 2.
    Neglected the part of the poloidal flux produced by the plasma current that lies outside of the plasma, which is most of the poloidal flux produced by the plasma current.
  3. 3.
    Assumed a spatially uniform perpendicular energy diffusivity to calculate the current profile that would be expected in a ramp down. Since that particular profile is stable from the 1987 study of TFTR disruptions by Cheng, Furth, and Boozer MHD stab , Fitzpatrick made the remarkable extrapolation that all ramp downs of similar speed are stable.
  4. 4.
    Gave a general criterion for avoiding disruptive states, which is clearly incorrect.
Fitzpatrick would have avoided most of these errors had he considered Figure 1 in my article Boozer:flux , and the relation between the poloidal flux enclosed by the magnetic axis of a tokamak and the loop voltage around that axis. As explained just below the figure, this relation follows from the application of Stokes’ Theorem to Faraday’s Law.
Fitzpatrick made no note of the sentence in the abstract of the submitted version of my article that gave the major implication on plasma shutdown: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.” Fitzpatrick did not attempt to determine what range of plasma current profiles are stable, and he omitted the majority of the poloidal flux produced by the plasma current, the part outside of the plasma.
Fitzpatrick’s certainty that my concerns about disruptions are unfounded relate to a statement in his September 10, 2025 email to me: “the average broadness of the current profile is easily controlled.” This statement by Fitzpatrick might make one think disruptions in tokamak would be rarely encountered and then only as a result of operator incompetence. It was not so simple in JET shutdowns Sozzi:2020 .
Section III discusses Fitzpatrick’s misrepresentation of my views. As stated by the first reviewer of my Physics of Plasmas submission Boozer:flux : “a more rigorous model was recently published in Nuclear Fusion by Fitzpatrick, prompted by the arXiv article from this author.” Fitzpatrick’s first reference was my arXiv article. His second reference was “Boozer A.H. 2025 private communication.” The “private communication” was a number of emails that also made my concern explicit—the smallness of the fractional poloidal flux change required to produce a disruptive state.
The extreme nature of the errors in science and of the misrepresentation of my views caused me to enquire, starting on November 24, 2025, whether Fitzpatrick’s paper is consistent with the scientific and ethical standards of Nuclear Fusion. On January 5, 2026 Nuclear Fusion said they had “no evidence of intentional misrepresentation,” despite the emails.
The urgency of having this Comment publicly available is explained in Section IV. Both the renewal of grants and the publication of a paper on quick assessments of issues faced by tokamak power plants and of strategies to circumvent them are dependent on the public availability of this Comment.
II Poloidal flux and its evolution in tokamaks
Figure 1: The lines of the poloidal magnetic field produced by the toroidal plasma current are shown together with the magnetic field produced by the central solenoid of a tokamak.
II.1 Basic properties the poloidal flux
Although ignored by Fitzpatrick, fundamental properties of the poloidal flux and its evolution were illustrated in Figure 1, from my article Boozer:flux , and a simple application of Stokes’ Theorem to Faraday’s Law at the circular magnetic axis of tokamak was given just below that figure.
The Faraday’s Law result is
$\frac{\partial}{\partial t}\int\vec{B}\cdot d\vec{a}$ $=$ $-\int\left(\vec{\nabla}\times\vec{E}\right)\cdot d\vec{a}$ (1)
$\frac{\partial\psi^{ax}_{p}}{\partial t}$ $=$ $V_{\ell}^{ax},$ (2)
The magnetic flux going down through the magnetic axis, which is a circle of radius $R_{ax}$ is
$\psi_{p}^{ax}\equiv-\int\vec{B}\cdot d\vec{a}$ (3)
integrated over the area enclosed by the axis. The loop voltage at the axis is
$V_{\ell}^{ax}$ $\equiv$ $\oint\vec{E}\cdot d\vec{\ell}$ (4)
$=$ $2\pi R_{ax}\eta_{ax}j_{ax}$ (5)
with $\eta_{ax}$ the resistivity at the axis and $j_{ax}$ the current density flowing along the magnetic field line that forms the axis.
The choice of sign for the poloidal flux in $(R,\varphi,Z)$ cylindrical coordinates,
$\psi_{p}(R)$ $\equiv$ $-2\pi\int_{0}^{R}B_{z}RdR,$ (6)
may seem odd, but it results in $\psi_{p}$ increasing with distance $r$ from the axis, which gives a positive safety factor. The distance from the axis is $r\equiv R_{ax}-R$, and $2\pi(R_{ax}-r)dr$ gives the area element for defining the poloidal flux that lies within a distance $r$ of the axis.
$\psi_{p}^{f^{\prime}}$ $\equiv$ $\psi_{p}(R)-\psi_{p}(R_{ax})$ (7)
$=$ $2\pi\int_{0}^{r}B_{z}(R_{ax}-r)dr,\hskip 14.45377pt\mbox{where}\hskip 14.45377pt$ (8)
$\simeq$ $2\pi R_{ax}\int_{0}^{r}B_{\theta}dr\hskip 14.45377pt\mbox{for $r<<R_{ax}$;}\hskip 14.45377pt$ (9)
The poloidal flux $\psi_{p}^{f^{\prime}}$ is related to Fitzpatrick’s definition in which he used a cylindrical $(r,\theta,z)$ model for the poloidal magnetic field in a tokamak.
Fitzpatrick’s definition of poloidal flux differed in two ways from $\psi_{p}^{f^{\prime}}$ of Equation 9. First, he included a factor of no fundamental importance due to the non-circularity of the magnetic surfaces, which for simplicity will be ignored here. The effect of magnetic-surface shape on the rotational transform is given by Equation (237) in Reference NF review . Second and what is of fundamental importance, Fitzpatrick did not include the poloidal flux that lies outside of the plasma but is produced by the plasma current; he retained only the poloidal flux internal to the plasma,
$\psi_{p}^{in}=2\pi R_{ax}\int_{0}^{a}B_{\theta}dr.$ (10)
He did include a time dependent value within his definition of the poloidal flux at the magnetic axis, Equation (14).
For the circular surfaces of a $(r,\theta,z)$ model, $B_{\theta}=\mu_{0}I(r)/(2\pi r)$, where $I(r)=2\pi\int_{0}^{r}j_{t}(r)rdr$ and $j_{t}$ is the net toroidal current density. Fitzpatrick set the contribution to $\psi_{p}^{f^{\prime}}$ for $r>a$ equal to zero, where $a$ is the plasma radius. In a cylindrical approximation, the poloidal flux that lies between $a$ and an outer radius $b>a$ is $\{\mu_{0}I_{p}/(2\pi)\}\int_{a}^{b}dr/r$, which is proportional to $\ln(b/a)$ and becomes logarithmically infinite for $b>>a$.
As noted in my article and stated in many other places, the poloidal flux of a large aspect ratio toroidal shell with major and minor radii $R$ and $a$ carrying a current $I_{p}$ is
$\psi_{p}^{ex}=\mu_{0}R\left\{\ln\left(\frac{8R}{a}\right)-2\right\}I_{p}.$ (11)
This expression can be obtained using the method of the solved problem 5.32 in Reference Jackson .
For standard tokamaks, the external part of the poloidal flux produced by the plasma current is larger than the flux given by the plasma current that lies within the plasma, the $\psi_{p}^{in}$ of Equation (10). In his Equation (21), Fitzpatrick gives the typical poloidal flux by his calculation as $\psi_{0}=2\pi R_{ax}aB_{\theta}(a)$. When one ignores the increased current required for a given flux due to non-circularity, which affects both $\psi_{p}^{in}$ and $\psi_{p}^{ex}$. Fitzpatrick’s typical flux is $\psi_{0}=\mu_{0}R_{ax}I_{p}$, and the ratio $\psi_{p}^{ex}/\psi_{0}=\ln(8R/a)-2\approx 1.178$ when $R/a=3$. For a parabolic current profile, the ratio of the external to the internal flux produced by the plasma current is even larger, $\psi_{p}^{ex}/\psi_{p}^{in}=1.57$.
The poloidal flux enclosed by the magnetic axis, Equation (3), has two important parts: the part produced by the plasma current $\Psi_{p}$ and the poloidal flux in the solenoid, $\psi_{sol}(t)$, which is the only directly controllable function of time;
$\psi_{p}^{ax}$ $=$ $\Psi_{p}+\psi_{sol}(t);$ (12)
$\Psi_{p}$ $=$ $-(\psi_{p}^{in}+\psi_{p}^{ex}).$ (13)
In the approximation of circular surfaces and a large aspect ratio, $R/a>>1$, Equation (10) gives $\psi_{p}^{in}$, and Equation (11) gives $\psi_{p}^{ex}$.
II.2 Disentangling Fitzpatrick’s expressions
Richard Fitzpatrick entangles his arguments about poloidal flux evolution with issues that are in principle separable, such as heat transport. This entanglement produces confusion—apparently to Fitzpatrick himself—and a disentanglement is required.
The first step in a disentanglement is determining what is precisely Fitzpatrick’s definition of the poloidal flux and what basic equations does it obey in the cylindrical $(r,\theta,z)$ approximation that he utilizes. His Equations (6) and (9) together with his Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the Fitzpatrick definition of the poloidal flux $\psi_{f}(r,t)$ is
$\psi_{f}(r,t)$ $=$ $\psi_{f}(0,t)+2\pi R_{ax}\int_{0}^{r}B_{\theta}dr;$ (14)
$\frac{d\psi_{f}(0,t)}{dt}$ $=$ $2\pi R_{ax}E_{z}$ (15)
$=$ $V_{\ell}^{ax},$ (16)
the loop voltage at the magnetic axis.
The poloidal magnetic flux is the flux enclosed in the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$. The time derivative of the poloidal flux satisfies exactly
$\frac{\partial\psi_{p}(\psi_{t},t)}{\partial t}$ $=$ $V_{\ell}(\psi_{t},t);$ (17)
$V_{\ell}(\psi_{t},t)$ $\equiv$ $\lim_{L\rightarrow\infty}\frac{\int_{-L}^{L}\vec{E}\cdot d\vec{\ell}}{\int_{-L}^{L}\vec{\nabla}\left(\frac{\varphi}{2\pi}\right)\cdot d\vec{\ell}}.$ (18)
This is true during any period that the magnetic surface exists, where $d\vec{\ell}$ is the differential distance along a magnetic field line in that surface. Unlike, the equivalent equation at the magnetic axis, where $\psi_{t}=0$, this is not trivially derived from Faraday’s Law and Stokes’ Theorem, but is given by Equations (27) and (31) in Reference Boozer:RMP .
A fundamental mathematical point that Fitzpatrick missed is that the toroidal magnetic flux is defined by an inside-outwards area integral from the magnetic axis while the poloidal flux is defined by an outward-inwards integral towards the magnetic axis.
Fitzpatrick’s condition for a sufficiently slow ramp up or down to avoid a disruption, his Equation (42), is equivalent to
$V_{\ell}^{ax}$ $>$ $\left|2\pi R_{ax}\frac{\partial}{\partial t}\int_{0}^{a}B_{\theta}dr\right|\hskip 14.45377pt\mbox{or}\hskip 14.45377pt$ (19)
$>$ $\left|\frac{d\psi_{p}^{in}}{dt}\right|.$ (20)
His condition can be expressed as a condition on the loop voltage at the plasma edge, $V_{\ell}^{edge}=(d\psi_{p}/dt)_{edge}$, which is at $r=a$. The poloidal flux at the plasma edge is given by
$\psi_{p}^{edge}(t)$ $=$ $\psi_{p}^{ax}-\psi_{p}^{in}\hskip 14.45377pt\mbox{or}\hskip 14.45377pt$ (21)
$=$ $\psi_{sol}(t)+\psi_{p}^{ex}.$ (22)
During a ramp up, $d\psi_{p}^{in}/dt>0$, Fitzpatrick’s condition, Equation (20), for avoiding a disruption is $V_{\ell}^{edge}>0$, which is a necessary condition for any ramp up but not a sufficient condition for avoiding a disruption. During a ramp-down, $d\psi_{p}^{in}/dt<0$, Fitzpatrick’s condition for avoiding a disruption is $V_{\ell}^{edge}<2V_{\ell}^{ax}$, which also not a sufficient condition for avoiding a disruption. What is required is that the current density profile not be disruptive as discussed in Subsection II.4.
II.3 Fitzpatrick check for non-disruptive profiles
Although Fitzpatrick gave an expression for the profile of perpendicular energy diffusivity of some generality in his Equation (88), he actually determined the current profile that would be obtained for only one case, a radially constant energy-diffusivity profile. He then checked that the resulting current profile is not disruptive using the internal inductance $\ell_{i}$, Fitzpatrick’s Equation (90), and the edge safety factor $q_{a}$ of the 1987 Cheng, Furth, and Boozer study MHD stab in a cylindrical approximation of disruptions in TFTR.
What Fitzpatrick did not do is determine the range of profiles of his energy diffusivity that give non-disruptive current profiles even by the Cheng et al criterion that he was using.
Showing stability for one profile of the heat diffusivity is certainly no guide for determining the danger of disruptions. It might be noted, that the general view among tokamak experts is that energy diffusivity is not the correct way to characterize energy transport. The temperature profiles are given rigidity by microturbulence. See for example Reference Holland:2021 . This rigidity is sometimes thought to ensure that a spatially constant loop voltage will give disruption-stable current profiles.
II.4 Actual condition for disruption avoidance
Not only did Fitzpatrick ignore the simple statements in my article about the poloidal flux and its evolution, he also ignored the statement in the abstract in my paper about the important result: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
Fitzpatrick could not examine the correctness of this statement for two reasons. First, by ignoring $\psi_{p}^{ex}$ he ignored most of the poloidal flux produced by the plasma current. Second, by not considering the range of profiles of current density that are consistent with stability by the Cheng-Furth-Boozer criterion MHD stab , which Fitzpatrick used in his paper, he could not estimate what range of poloidal fluxes are stable to disruptions.
Only a small fractional change in the poloidal flux is required to move a tokamak equilibrium that is stable to disruptions into a disruptive state. Consequently, a very careful control of the poloidal flux evolution is required during a plasma shut down—all the poloidal flux produced by the plasma current $\Psi_{p}=-(\psi_{p}^{in}+\psi_{p}^{ex})$ must be removed. The length of the timescale for $\Psi_{p}$ removal is so long that the current profile within the plasma has time to relax to nearly its steady-state limit, which is a spatially constant loop voltage $V_{\ell}$ across the plasma. Ignoring current drive and the bootstrap current, as Fitzpatrick does, the net toroidal current density on each magnetic surface, which is denoted by its enclosed toroidal flux $\psi_{t}$, is
$j_{t}(\psi_{t})=\frac{V_{\ell}}{2\pi R\eta(Z_{eff},T_{e})},$ (23)
where $2\pi R$ is the average distance required for a field line to make a toroidal circuit, $Z_{eff}$ is the correction to the resistivity produced by impurities, and $T_{e}$ is the electron temperature. Neoclassical effects to the resistivity $\eta$ can be easily included. The difficulty of obtaining disruption-free shutdowns in JET was discussed in Sozzi:2020 . This was thought largely due to impurity issues.
In pulsed tokamaks power plants, the pulse length is set by the length of time the central solenoid can supply the loop voltage required to maintain the current. Typically the solenoid is designed to supply a current-maintenance flux comparable to $\Psi_{p}$ during the fusion-production flattop, which implies the current profile of Equation (23) with $V_{\ell}$ essentially the same on all magnetic surfaces in the plasma.
The actual condition for disruption avoidance is that the current profile given by Equation (23), with $V_{\ell}$ the same on all magnetic surfaces, never crosses into a profile unstable to a disruption. This crossing can be estimated using Cheng et al MHD stab , but a far more reliable estimate could be obtained using three-dimensional magnetohydrodynamics codes coupled with comparisons to experiments.
Although not emphasized in his paper, Fitzpatrick assumed a spatially constant loop voltage in the plasma region in both ramp-up and ramp-down of the plasma current. The primary point in my Physics of Plasmas submission was the spatial constancy of the loop voltage. A major difference in calculations is that my article considered a number of current profiles that ranged over the stable region of the Cheng et al MHD stab paper and found the poloidal flux in the plasma changed by relatively small amount over that range of current profiles. Fitzpatrick considered only one current profile, which happened to be stable, but this is not a credible argument for the robust stability against disruptions for all realistic current profiles.
III Fitzpatrick’s misrepresentations
Fitzpatrick made it clear in the Abstract, Introduction, and in the Summary and Discussion that he wrote his paper Fitzpatrick:2026 to correct an arXiv article of mine, his first reference. That would be fine except he was misinterpreting my article. His second reference was “Boozer A.H. 2025 private communication.”
The “private communication” was a chain of emails some even before Richard submitted his article to Nuclear Fusion on August 18. These emails politely but explicitly stated that he was misinterpreting me, that I was revising my article to remove any confusion, and what I viewed as the most important point of my article. If Fitzpatrick had just ignored these emails, that would have been irresponsible. But, Richard coupled his first and second references everywhere except in the Abstract to reinforce his claim to be correcting a false view of mine.
His misrepresentation of “private communication” is particularly clear in an August 13 email exchange because it was before his paper had been submitted to Nuclear Fusion. My email to Fitzpatrick said “Actually I didn’t really mean the expected ramp down time was 1000 sec, and it would take a lot of power to keep the plasma going. What I was trying to indicate is that a large cooling of the plasma is necessary but the profile had to remain relatively fixed while doing that ramp down. I agree my words were confusing, and they have been changed.” Fitzpatrick responded: “Would it be accurate to say that you think that the ramp-down times are longer than the SPARC and ITER people are claiming” My response: “The primary issues are the energy and magnetic flux input into the plasma that is required during a shutdown to ensure the current profile does not slip into a disruptive state more than one time in at least a thousand shut-downs. A major problem in my view, is that a 20% change in the poloidal flux distribution is sufficient to take a plasma far from a state as far as possible from a disruption into a disruption. This means a disruption can occur in approximately a fifth of the time one would expect a flattop to last.”
The perception of the first reviewer of my Physics of Plasmas submission Boozer:flux is revealing: “a more rigorous model was recently published in Nuclear Fusion by Fitzpatrick, prompted by the arXiv article from this author.” Fitzpatrick told me that he was not the reviewer in a response to a version of this Comment that I sent to him for comments. The first reviewer perceived that the whole motivation for Fitzpatrick’s paper Fitzpatrick:2026 was my arXiv article but did not consider the statement of purpose in the abstract of that article.
The last “private communication” that Fitzpatrick and I had was on 9-10 September 2025, which was before he submitted his revised version of his article, 25 September 2025. This was the version that was accepted. Since this communication illustrates our different perceptions of physics issues, it is quoted exactly below. In his article, Fitzpatrick did not ascribe to me the views that I stated, nor does he address in this email or in his paper his confusion about the definition and the evolution of the poloidal flux.
Boozer-Fitzpatrick 9-10 September emails
From: Richard Fitzpatrick rfitzp@utexas.edu Subject: Date: Re: Revised abstract & introduction to “Constraints on the magnetic field evolution in tokamak power plants” September 10, 2025 at 4:27 PM To: Allen H. Boozer ahb17@columbia.edu
Dear Allen I do not really agree that tokamak plasma are “self-organized”. When I think of the term self-organized, I naturally think of an RFP plasma. Such plasmas are subject to continuous large-amplitude m=1 and m=0 tearing modes. It is these modes that relax the plasma toward a Taylor state. However, a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central q value from falling significantly below unity. Given that q(0) is pinned to a value close to unity, and q(a) is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities. Richard
On Tue, Sep 9, 2025 at 4:29 PM Allen H. Boozer <<ahb17@columbia.edu>> wrote: The confusions on the important points in my arXiv article “Constraints on the magnetic field evolution in tokamak power plants” were convincing that the abstract and introduction had to be rewritten. Any comments are appreciated, especially ones on balance and fair quoting. Discussions by Zoom or phone are always welcome, Allen 646-319-9962
A copy of my article Boozer:flux was attached to my email and a plaintext version of its abstract was included.
IV Urgency
Two issues make the publication of this Comment urgent.
  1. 1.
    The U.S. Department of Energy rejected the renewal proposal that would have provided support for this paper on December 1, 2025 after it had been under review for twenty-three months. This was shortly after Fitzpatrick’s Nuclear Fusion paper appeared online, November 6, 2025. It is not only important that my funding for this type of work be restored, but also one other renewal proposal be funded, which has been under review for more than eight months, has not yet had a response.
    The renewal of proposals based on a continuing development of fundamentally new concepts and inventions has become difficult under the new Fusion Science and Technology Roadmap DoE:2025 . Page 7 of the Roadmap prominently displays: “Innovate and advance the science and engineering of fusion with well-defined milestones and metrics.” As interpreted, “well-defined milestones” is consistent with funding based on the development of fundamentally new concepts and inventions only when their timing can be predicted years into the future—a logical impossibility.
    Having my contributions trashed in a respected journal, even in an article filled with scientific errors and misinterpretations does not help. It is natural to assume that Nuclear Fusion, a respected journal, would not publish anything with so many problems—and would not stand behind such a paper if it were published.
  2. 2.
    The Physics of Plasmas has said that they would consider a new submission of my article Constraints on the magnetic field evolution in tokamak power plants only after the incorrectness of the charges in Richard Fitzpatrick’s paper against my article had been demonstrated.
    My article allows quick assessments of issues faced by tokamak power plants and of strategies to circumvent them. At present a number of privately funded proposals for tokamak power plants are being developed, and such assessments are of great practical importance. There is no reason that a paper filled with scientific mistakes and misrepresentations should be allowed to continue to block a publication important to power-plant proposals.