### v2

# A Simple Model of Current Ramp-Up and Ramp-Down in Tokamaks

Abstract
[…]
I Introduction
Richard Fitzpatrick’s 2026 article Fitzpatrick:2026 on current ramp-up and ramp-down in tokamaks, confusingly coupled a number of effects that are in principle separable. A disentanglement, Section II, shows his fundamental physics errors in the definition and evolution of the poloidal flux in tokamaks. In particular, Fitzpatrick:
  1. 1.
    Did not explicitly include the changes in the flux in the central solenoid, $\psi_{sol}$, in his analysis, which is the only flux that is a directly controllable function of time.
  2. 2.
    Neglected the part of the poloidal flux produced by the plasma current that lies outside of the plasma, which is most of the poloidal flux produced by the plasma current.
  3. 3.
    Assumed a spatially uniform perpendicular energy diffusivity to calculate the current profile that would be expected in a ramp down. Since that particular profile is stable from the 1987 study of TFTR disruptions by Cheng, Furth, and Boozer MHD stab , Fitzpatrick made the remarkable extrapolation that all ramp downs of similar speed are stable.
  4. 4.
    Gave a general criterion for avoiding disruptive states, which is clearly incorrect.
Fitzpatrick would have avoided most of these errors had he considered Figure 1 in my article Boozer:flux , and the relation between the poloidal flux enclosed by the magnetic axis of a tokamak and the loop voltage around that axis. As explained just below the figure, this relation follows from the application of Stokes’ Theorem to Faraday’s Law.
Figure 1: The lines of the poloidal magnetic field produced by the toroidal plasma current are shown together with the magnetic field produced by the central solenoid of a tokamak.
Fitzpatrick has insisted on misinterpreting my article as primarily about acceptable ramp-up and ramp-down times. The abstract of his paper states: “there is no indication that the design ramp times are infeasible, as was recently suggested in Boozer (2025 arXiv:2507.05456v1).” Somehow he came to this conclusion when even the first arXiv version, which he emphasizes rather than the version submitted to the Physics of Plasmas, stated the long decay time of the plasma current at fusion temperatures means: “A rapid cooling of the central part of the plasma seems required, but unless this is done with care, the current profile will evolve into a disruptive state.” The same arXiv version also said the restart period in a pulsed tokamak has significant freedom: “In principle, the restart period can be made arbitrarily short by clever choices of the time dependence of $\psi_{sol}(t)$ and the heating power. Of course, the periods in which substantial heating power is required must be short compared to the periods of fusion burn to have net energy to sell.” Fitzpatrick made no note of the sentence in the abstract of the submitted version of my article: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
As discussed in Section II.4, the smallness of the required change in the poloidal flux to cross the full range of stable plasma current profiles implies the current profile in the plasma is determined by the spatial constancy of the loop voltage. Actually Fitzpatrick assumed a constant loop voltage to obtain the single profile of plasma current that he discussed in his Nuclear Fusion paper, but did not explain why this was an accurate approximation.
Despite a number of emails that attempted to correct Fitzpatrick’s mistaken interpretation of my article, he not only persisted in that misinterpretation but reinforced his misinterpretation in every reference to my arXiv article, except in the abstract, by citing “private communication” as well. The only private communication that we had was the emails that tried to correct his misinterpretation.
The reason for Fitzpatrick’s persistence on misinterpreting my article can be explained by his statement of dogma in his September 10, 2025 email, part of his “private communication” reference: “a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central $q$ value from falling significantly below unity. Given that $q(0)$ is pinned to a value close to unity, and $q(a)$ is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities.”
Fitzpatrick’s dogma seems to imply that disruptions could be easily avoided by a competent machine operator. It was not so simple in JET shutdowns Sozzi:2020 .
Although I have never said that tokamaks have “an inevitable drift to a disruptive state,” ensuring tokamaks have a sufficiently low disruption rate for the economic feasibility of power plants is non-trivial and requires improved understanding.
A similar dogma to the one that prevented Fitzpatrick from acknowledging my views seems to affect the judgement of a significant part of the tokamak community and allows them to overlook even obvious mistakes in science. To correct this, the scientific evidence that does not conform with Fitzpatrick’s statement of dogma needs to be given.
A Physics of Plasmas reviewer of my article said Fitzpatrick’s paper is “a more rigorous model.” A leader in the tokamak program stated that Fitzpatrick’s paper: “refutes your arguments that this (disruptions) inherently dooms tokamaks. This reason goes well beyond Richard’s paper, which attacks the problem from a fundamental and transparent way in a simple model. This is because some of the most sophisticated and benchmarked numerical simulations we have for tokamaks are time-dependent stability/equilibrium codes that tackle the inductive problem. These have been developed over decades of development and testing against thousands of tokamaks discharges on an extremely large range of device sizes, plasma temperatures and engineering design.” Similar comments have come from a number of persons.
It is surprising that anyone believes the quality of the science underlying the design of tokamak power plants is at such a low level that Fitzpatrick’s paper adds important information about disruption avoidance. The glowing descriptions of “sophisticated and benchmarked numerical simulations we have for tokamaks” is an appropriate statement to make to a potential private funder but is not a scientific response to a scientific study into the requirements for avoiding disruptions.
In addition to theory and simulations, empirical studies advance understanding of disruptions and their avoidance. Studies based on the DCAF code developed by Steve Sabbagh’s group have been used to determine Sabbagh:2024 the empirical dependence of tokamak disruptions on plasma quantities such as the internal inductance $\ell_{i}$ and edge safety factor $q_{95}$, Figure 2, The results for MAST-U, which has highly non-circular magnetic surfaces, give preliminary evidence that an internal inductance that is either too low or too high gives disruptive plasmas with a stable region between. An empirical study that is related to that of Figure 2 is given in Figure 3, which is a study of disruptions in JET with an ITER-like wall, JET-dis:2020 . That figure shows that internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ are not sufficient to define a stable region. Are the “sophisticated and benchmarked numerical simulations we have for tokamaks” adequate for explaining Figures 2 and 3?
Figure 2: MHD-caused disruptions are illustrated in an internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ diagram with separate symbols for KSTAR and NSTX-U as well as for two operational years of MAST-U. Two branches of DCS trigger points can be recognized in both MAST-U experimental years. The two branches (visualized with green lines) are separated along the $\ell_{i}$ axis into two regions, with the internal inductance being roughly restricted to $1.5\lesssim\ell_{i}\lesssim 1.7$ in branch 1 and $0.8\lesssim\ell_{i}\lesssim 1.2$ in branch 2 . The region between the branches is free of MHD-caused disruptions for $q_{95}\gtrsim 6.3.$ This is Figure 7 in the paper V. Zamkovska, S.A. Sabbagh, M. Tobin, et al, Nucl. Fusion 64, 066030 (2024).
Figure 3: Pre-disruptive parameters in JET with an ITER-like wall are shown in a $\ell_{i}-q_{95}$ stability diagram. This was Figure 13 in Gerasimov et al, Nucl. Fusion 60, 066028 (2020). As the authors note: “It may be expected that a disruption free space may be defined in the $\ell_{i}-q_{95}$ empirical stability diagram, assuming that plasma current profiles tend to maintain itself inside the permissible values. In reality, the JET-ILW pre-disruptive plasma equilibrium parameters create a diffused cloud on the $\ell_{i}-q_{95}$ stability diagram without room for non-disruptive plasmas.”
Ensuring disruptions are sufficiently rare for the feasibility of tokamak power plants would seem to require identifying the externally controllable parameters that can be localized in a way that makes the plasma safe against disruptions. This requirement is more stringent than in existing tokamaks or in JET because of the limitations on diagnostics and actuators in power plants. Generative AI using data from both experiments and simulations could help in this identification.
A major determinant of disruptivity is the profile of the net plasma current $I(\psi_{t})$ enclosed by a magnetic surface that also encloses a toroidal magnetic flux $\psi_{t}$. The derivative $I^{\prime}(\psi_{t})\equiv dI/d\psi_{t})$ is given by the divergence-free part of $\vec{j}_{||}=(j_{||}/B)\vec{B}$, which is the net plasma current, and is determined by the spatial constancy of the loop voltage. Despite the dogma expressed by Fitzpatrick and apparently accepted by many in the tokamak community, the naturally arising profiles of $I(\psi_{t})$ during startup, flattop, and shutdown have many uncertainties. Figures 2 and 3 show that actual tokamak plasmas have many values of $\ell_{i}$, which quantifies the peakedness of the current profile, for a given edge safety factor. There is not just the one current profile that Fitzpatrick focused on.
What theory can do and what I am doing is identifying controllable features in $I(\psi_{t})$ that determine the internal inductance $\ell_{i}$. When the central safety factor $q_{0}$ is held to unity by sawteeth, the derivative $I^{\prime}(0)$ is defined. The determination of the current density $I^{\prime}(\psi_{t})$ by the loop voltage implies $I^{\prime}(\psi_{t})\geq 0$ without negative external current drive. Both a positive $I^{\prime}$ near the edge, as in a H-mode plasma, and triangular shaping of the magnetic surfaces reduce $\ell_{i}$ for a fixed edge to central $q$-factor ratio with an otherwise fixed $I(\psi_{t})$. It would be interesting to determine if extending the empirical data used to construct Figures 2 and 3 would be much more determinate using a four than in a two parameter fit.
The acceptance of a Fitzpatrick-like dogma apparently led to strong negative reactions to two statements that were in the various versions of my arXiv article: The first statement was: “Diversity of concepts is important for the development and optimization of fusion. Nevertheless, an intelligent allocation of resources among the various concepts requires theoretical and computational assessments of the risks, time, and costs of fusion development based on these concepts.” In principle, it would seem anyone who believes in the importance of the rapid development of fusion would agree. The second statement was: “The purpose of this paper is not to remove tokamaks from consideration but rather to determine on what issues research should be focused for tokamaks to be a credible alternative to stellarators. For this, the obvious strategy is to determine what actuators are available to control the things that must be controlled and what plasma data could be available to guide these actuators.” This statement seems controversial only when one believes the external-current determination of stellarator plasmas is not fundamentally different in disruption and other plasma control issues when compared to the effects of internal plasma currents in tokamaks.
Anyone responsible for managing the development of fusion power plants would do well to ponder Admiral Rickover’s well-known quotes based on his success in the launching the first nuclear submarine about fifteen years after the splitting of the uranium nucleus in Berlin. An important quote is: “It is a human inclination to hope things will work out, despite evidence or doubt to the contrary. A successful manager must resist this temptation.”
II Poloidal flux and its evolution in tokamaks
II.1 Basic properties the poloidal flux
Although ignored by Fitzpatrick, fundamental properties of the poloidal flux and its evolution were illustrated in Figure 1, from my article Boozer:flux , and a simple application of Stokes’ Theorem to Faraday’s Law at the circular magnetic axis of tokamak was given just below that figure.
The Faraday’s Law result is
$\frac{\partial}{\partial t}\int\vec{B}\cdot d\vec{a}$ $=$ $-\int\left(\vec{\nabla}\times\vec{E}\right)\cdot d\vec{a}$ (1)
$\frac{\partial\psi^{ax}_{p}}{\partial t}$ $=$ $V_{\ell}^{ax},$ (2)
The magnetic flux going down through the magnetic axis, which is a circle of radius $R_{ax}$ is
$\psi_{p}^{ax}\equiv-\int\vec{B}\cdot d\vec{a}$ (3)
integrated over the area enclosed by the axis. The loop voltage at the axis is
$V_{\ell}^{ax}$ $\equiv$ $\oint\vec{E}\cdot d\vec{\ell}$ (4)
$=$ $2\pi R_{ax}\eta_{ax}j_{ax}$ (5)
with $\eta_{ax}$ the resistivity at the axis and $j_{ax}$ the current density flowing along the magnetic field line that forms the axis.
The choice of sign for the poloidal flux in $(R,\varphi,Z)$ cylindrical coordinates,
$\psi_{p}(R)$ $\equiv$ $-2\pi\int_{0}^{R}B_{z}RdR,$ (6)
with $Z=0$, may seem odd, but it results in $\psi_{p}$ increasing with distance $r$ from the magnetic axis, which gives a positive safety factor. The distance from the axis is $r\equiv R_{ax}-R$, and $2\pi(R_{ax}-r)dr$ gives the area element for defining the poloidal flux that lies within a distance $r$ of the axis.
$\psi_{p}^{f^{\prime}}$ $\equiv$ $\psi_{p}(R)-\psi_{p}(R_{ax})$ (7)
$=$ $2\pi\int_{0}^{r}B_{z}(R_{ax}-r)dr,\hskip 14.45377pt\mbox{where}\hskip 14.45377pt$ (8)
$\simeq$ $2\pi R_{ax}\int_{0}^{r}B_{\theta}dr\hskip 14.45377pt\mbox{for $r<<R_{ax}$;}\hskip 14.45377pt$ (9)
The poloidal flux $\psi_{p}^{f^{\prime}}$ is related to Fitzpatrick’s definition in which he used a cylindrical $(r,\theta,z)$ model for the poloidal magnetic field in a tokamak.
Fitzpatrick’s definition of poloidal flux differed in two ways from $\psi_{p}^{f^{\prime}}$ of Equation 9. First, he included a factor of no fundamental importance due to the non-circularity of the magnetic surfaces, which for simplicity will be ignored here. The effect of magnetic-surface shape on the rotational transform is given by Equation (237) in Reference NF review . Second and what is of fundamental importance, Fitzpatrick did not include the poloidal flux that lies outside of the plasma but is produced by the plasma current; he retained only the poloidal flux internal to the plasma,
$\psi_{p}^{in}=2\pi R_{ax}\int_{0}^{a}B_{\theta}dr.$ (10)
He did include a time dependent value within his definition of the poloidal flux at the magnetic axis, Equation (14).
For the circular surfaces of a $(r,\theta,z)$ model, $B_{\theta}=\mu_{0}I(r)/(2\pi r)$, where $I(r)=2\pi\int_{0}^{r}j_{t}(r)rdr$ and $j_{t}$ is the net toroidal current density. Fitzpatrick set the contribution to $\psi_{p}^{f^{\prime}}$ for $r>a$ equal to zero, where $a$ is the plasma radius. In a cylindrical approximation, the poloidal flux that lies between $a$ and an outer radius $b>a$ is $\{\mu_{0}I_{p}/(2\pi)\}\int_{a}^{b}dr/r$, which is proportional to $\ln(b/a)$ and becomes logarithmically infinite for $b>>a$.
As noted in my article and stated in many other places, the poloidal flux of a large aspect ratio toroidal shell with major and minor radii $R$ and $a$ carrying a current $I_{p}$ is
$\psi_{p}^{ex}=\mu_{0}R\left\{\ln\left(\frac{8R}{a}\right)-2\right\}I_{p}.$ (11)
This expression can be obtained using the method of the solved problem 5.32 in Reference Jackson .
For standard tokamaks, the external part of the poloidal flux produced by the plasma current is larger than the flux given by the plasma current that lies within the plasma, the $\psi_{p}^{in}$ of Equation (10). In his Equation (21), Fitzpatrick gives the typical poloidal flux by his calculation as $\psi_{0}=2\pi R_{ax}aB_{\theta}(a)$. When one ignores the increased current required for a given flux due to non-circularity, which affects both $\psi_{p}^{in}$ and $\psi_{p}^{ex}$. Fitzpatrick’s typical flux is $\psi_{0}=\mu_{0}R_{ax}I_{p}$, and the ratio $\psi_{p}^{ex}/\psi_{0}=\ln(8R/a)-2\approx 1.178$ when $R/a=3$. For a parabolic current profile, the ratio of the external to the internal flux produced by the plasma current is even larger, $\psi_{p}^{ex}/\psi_{p}^{in}=1.57$.
The poloidal flux enclosed by the magnetic axis, Equation (3), has two important parts: the part produced by the plasma current $\Psi_{p}$ and the poloidal flux in the solenoid, $\psi_{sol}(t)$, which is the only directly controllable function of time;
$\psi_{p}^{ax}$ $=$ $\Psi_{p}+\psi_{sol}(t);$ (12)
$\Psi_{p}$ $=$ $-(\psi_{p}^{in}+\psi_{p}^{ex}).$ (13)
In the approximation of circular surfaces and a large aspect ratio, $R/a>>1$, Equation (10) gives $\psi_{p}^{in}$, and Equation (11) gives $\psi_{p}^{ex}$.
II.2 Disentangling Fitzpatrick’s expressions
Richard Fitzpatrick entangles his arguments about poloidal flux evolution with issues that are in principle separable, such as heat transport. This entanglement produces confusion—apparently to Fitzpatrick himself—and a disentanglement is required.
The first step in a disentanglement is determining what is precisely Fitzpatrick’s definition of the poloidal flux and what basic equations it obeys in the cylindrical $(r,\theta,z)$ approximation that he utilizes. His Equations (6) and (9) together with his Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the Fitzpatrick definition of the poloidal flux $\psi_{f}(r,t)$ is
$\psi_{f}(r,t)$ $=$ $\psi_{f}(0,t)+2\pi R_{ax}\int_{0}^{r}B_{\theta}dr;$ (14)
$\frac{d\psi_{f}(0,t)}{dt}$ $=$ $2\pi R_{ax}E_{z}$ (15)
$=$ $V_{\ell}^{ax},$ (16)
the loop voltage at the magnetic axis.
The poloidal magnetic flux is the flux enclosed in the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$.
The time derivative of the poloidal flux outside a magnetic surface that encloses toroidal flux $\psi_{t}$ satisfies exactly
$\frac{\partial\psi_{p}(\psi_{t},t)}{\partial t}$ $=$ $V_{\ell}(\psi_{t},t);$ (17)
$V_{\ell}(\psi_{t},t)$ $\equiv$ $\lim_{L\rightarrow\infty}\frac{\int_{-L}^{L}\vec{E}\cdot d\vec{\ell}}{\int_{-L}^{L}\vec{\nabla}\left(\frac{\varphi}{2\pi}\right)\cdot d\vec{\ell}}.$ (18)
This is true during any period that the magnetic surface exists, where $d\vec{\ell}$ is the differential distance along a magnetic field line in that surface. Unlike, the equivalent equation at the magnetic axis, where $\psi_{t}=0$, this is not trivially derived from Faraday’s Law and Stokes’ Theorem, but is given by Equations (27) and (31) in Reference Boozer:RMP .
A fundamental mathematical point that Fitzpatrick missed is that the toroidal magnetic flux is defined by an inside-outwards area integral from the magnetic axis while the poloidal flux is defined by an outward-inwards integral towards the magnetic axis.
Fitzpatrick’s condition for a sufficiently slow ramp up or down to avoid a disruption, his Equation (42), is equivalent to
$V_{\ell}^{ax}$ $>$ $\left|2\pi R_{ax}\frac{\partial}{\partial t}\int_{0}^{a}B_{\theta}dr\right|\hskip 14.45377pt\mbox{or}\hskip 14.45377pt$ (19)
$>$ $\left|\frac{d\psi_{p}^{in}}{dt}\right|.$ (20)
His condition can be expressed as a condition on the loop voltage at the plasma edge, $V_{\ell}^{edge}=(d\psi_{p}/dt)_{edge}$, which is at $r=a$. The poloidal flux at the plasma edge is given by
$\psi_{p}^{edge}(t)$ $=$ $\psi_{p}^{ax}-\psi_{p}^{in}\hskip 14.45377pt\mbox{or}\hskip 14.45377pt$ (21)
$=$ $\psi_{sol}(t)+\psi_{p}^{ex}.$ (22)
During a ramp up, $d\psi_{p}^{in}/dt>0$, Fitzpatrick’s condition, Equation (20), for avoiding a disruption is $V_{\ell}^{edge}>0$, which is a necessary condition for any ramp up but not a sufficient condition for avoiding a disruption. During a ramp-down, $d\psi_{p}^{in}/dt<0$, Fitzpatrick’s condition for avoiding a disruption is $V_{\ell}^{edge}<2V_{\ell}^{ax}$, which also not a sufficient condition for avoiding a disruption. What is required is that the current density profile not be disruptive as discussed in Subsection II.4.
II.3 Fitzpatrick check for non-disruptive profiles
Although Fitzpatrick gave an expression for the profile of perpendicular energy diffusivity of some generality in his Equation (88), he actually determined the current profile that would be obtained for only one case, a radially constant energy-diffusivity profile. He then checked that the resulting current profile is not disruptive using the internal inductance $\ell_{i}$, Fitzpatrick’s Equation (90), and the edge safety factor $q_{a}$ of the 1987 Cheng, Furth, and Boozer study MHD stab in a cylindrical approximation of disruptions in TFTR.
What Fitzpatrick did not do is determine the range of profiles of his energy diffusivity that give non-disruptive current profiles even by the Cheng et al criterion that he was using. Showing stability for one profile of the heat diffusivity is certainly no guide for determining the danger of disruptions.
In his response to an earlier version of this Comment, Fitzpatrick stated Fitzpatrick:2026A : “The transport of electron energy in tokamak plasma has been the subject of extensive research for the last 50 years. There is overwhelming evidence that such transport is diffusive in nature.”
Despite the dogmatism of Fitzpatrick’s statement, the general view among tokamak experts is that energy diffusivity is not always the correct way to characterize energy transport. The temperature profiles can be given rigidity by microturbulence, Holland:2021 ; Turb-profiles:2025 . The study by N.T. Howard et al of microturbulent transport in ITER Turb-profiles:2025 has sufficient information to calculate the loop voltage and determine its spatial constancy. This would determine the consistency of their assumed current profile $I(\psi_{t})$. Much work remains to establish the consistency of current profiles with both thermal transport and stability against disruptions in full ramp-up, flattop, and ramp-down scenarios even as the nuclear power production turns on and off.
II.4 Actual condition for disruption avoidance
Not only did Fitzpatrick ignore the simple statements in my article about the poloidal flux and its evolution, he also ignored the statement in the abstract in my paper about the important result: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
Fitzpatrick could not examine the correctness of this statement for two reasons. First, by ignoring $\psi_{p}^{ex}$ he ignored most of the poloidal flux produced by the plasma current. Second, by not considering the range of profiles of current density that are consistent with stability by the Cheng-Furth-Boozer criterion MHD stab , which Fitzpatrick used in his paper, he could not estimate what range of poloidal fluxes are stable to disruptions.
Only a small fractional change in the poloidal flux is required to move a tokamak equilibrium that is stable to disruptions into a disruptive state. Consequently, a very careful control of the poloidal flux evolution is required during a plasma shut down—all the poloidal flux produced by the plasma current $\Psi_{p}=-(\psi_{p}^{in}+\psi_{p}^{ex})$ must be removed. The length of the timescale for $\Psi_{p}$ removal is so long that the current profile within the plasma has time to relax to nearly its steady-state limit, which is a spatially constant loop voltage $V_{\ell}$ across the plasma. Ignoring current drive and the bootstrap current, as Fitzpatrick does, the net toroidal current density on each magnetic surface, which is denoted by its enclosed toroidal flux $\psi_{t}$, is
$j_{t}(\psi_{t})=\frac{V_{\ell}}{2\pi R\eta(Z_{eff},T_{e})},$ (23)
where $2\pi R$ is the average distance required for a field line to make a toroidal circuit, $Z_{eff}$ is the correction to the resistivity produced by impurities, and $T_{e}$ is the electron temperature. Neoclassical effects to the resistivity $\eta$ can be easily included. The difficulty of obtaining disruption-free shutdowns in JET was discussed in Sozzi:2020 . This was thought largely due to impurity issues.
In pulsed tokamaks power plants, the pulse length is set by the length of time the central solenoid can supply the loop voltage required to maintain the current. Typically the solenoid is designed to supply a current-maintenance flux comparable to $\Psi_{p}$ during the fusion-production flattop, which implies the current profile of Equation (23) with $V_{\ell}$ essentially the same on all magnetic surfaces in the plasma.
The actual condition for disruption avoidance is that the current profile given by Equation (23), with $V_{\ell}$ the same on all magnetic surfaces, never crosses into a profile unstable to a disruption. This crossing can be estimated using Cheng et al MHD stab , but a far more reliable estimate could be obtained using three-dimensional magnetohydrodynamics codes coupled with comparisons to experiments.
Although not emphasized in his paper, Fitzpatrick assumed a spatially constant loop voltage in the plasma region in both ramp-up and ramp-down of the plasma current. The primary point in my Physics of Plasmas submission was the spatial constancy of the loop voltage. A major difference in calculations is that my article considered a number of current profiles that ranged over the stable region of the Cheng et al MHD stab paper and found the poloidal flux in the plasma changed by relatively small amount over that range of current profiles. Fitzpatrick considered only one current profile, which happened to be stable, but this is not a credible argument for the robust stability against disruptions for all realistic current profiles.