### v3

# A Simple Model of Current Ramp-Up and Ramp-Down in Tokamaks

Abstract
[…]
I Introduction
Richard Fitzpatrick’s 2026 Nuclear Fusion paper Fitzpatrick:2026 on current ramp-up and ramp-down in tokamaks, confusingly coupled a number of effects that are in principle separable. A disentanglement, Section II, shows the paper contains fundamental physics errors in the definition and evolution of the poloidal flux in tokamaks. In particular, that paper Fitzpatrick:2026 :
  1. 1.
    Did not explicitly include the changes in the flux in the central solenoid, $\psi_{sol}$, which is the only flux that is a directly controllable function of time.
  2. 2.
    Neglected the part of the poloidal flux produced by the plasma current that lies outside of the plasma, which is most of the poloidal flux produced by the plasma current.
  3. 3.
    Assumed a time-independent profile profile, which was a spatial constant, of the thermal energy diffusivity to calculate the current profile that would be expected in a ramp down. Since that particular profile is stable from the 1987 study of TFTR disruptions by Cheng, Furth, and Boozer MHD stab , a remarkable extrapolation was made that all ramp downs of similar speed are stable.
  4. 4.
    Assumed that the profile of the poloidal magnetic field was essentially independent of time. Constancy would imply that if the initial current profile were stable to disruptions, it would remain stable.
The Nuclear Fusion paper Fitzpatrick:2026 was inspired by an arXiv article by Allen Boozer Boozer:flux and would have avoided most of these errors had Figure 1, which was in the arXiv article, been considered. This figure illustrates the relation between the poloidal flux enclosed by the magnetic axis of a tokamak and the loop voltage around that axis. As explained in the arXiv article Boozer:flux just below the figure, this relation follows from the application of Stokes’ Theorem to Faraday’s Law.
Figure 1: The lines of the poloidal magnetic field produced by the toroidal plasma current are shown together with the magnetic field produced by the central solenoid of a tokamak.
The abstract to the Nuclear Fusion paper Fitzpatrick:2026 states: “there is no indication that the design ramp times are infeasible, as was recently suggested in Boozer (2025 arXiv:2507.05456v1).” Even the first version of the arXiv article Boozer:flux , which is the focus of the paper Fitzpatrick:2026 , stated the long decay time of the plasma current at fusion temperatures means: “A rapid cooling of the central part of the plasma seems required, but unless this is done with care, the current profile will evolve into a disruptive state.” The same arXiv version also said the restart period in a pulsed tokamak has significant freedom: “In principle, the restart period can be made arbitrarily short by clever choices of the time dependence of $\psi_{sol}(t)$ and the heating power. Of course, the periods in which substantial heating power is required must be short compared to the periods of fusion burn to have net energy to sell.” The Nuclear Fusion paper Fitzpatrick:2026 makes no note of the sentence in the abstract of the submitted September 9, 2025 version of the arXiv article Boozer:flux : “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
As discussed in Section II.4, the smallness of the required change in the poloidal flux to cross the full range of stable plasma current profiles implies that current profiles in the plasma, when stable, are determined by the spatial constancy of the loop voltage. Just before Equation (42) of Fitzpatrick:2026 , an even stronger assumption was made. It was assumed that the profile of the poloidal field remains essentially fixed in time, which implies a spatially constant loop voltage. A constant poloidal-field profile ensures that if the original current profile is disruption stable, then it will remain stable. However, a spatially constant loop voltage does not imply the profile of the poloidal field remains essentially fixed in time since the resistivity, on which the loop voltage depends, is a function of the electron temperature $T_{e}$ and the $Z_{eff}$ due to impurities. The profiles of both $T_{e}$ and $Z_{eff}$ are generally time dependent.
Despite a number of emails that attempted to correct the mistaken interpretation of the arXiv article Boozer:flux , before the Nuclear Fusion paper Fitzpatrick:2026 was published, the published version reinforced that misinterpretation by the citations “private communication” in every reference to the arXiv article Boozer:flux , except in the abstract. Each “private communication” was an email. An insight into the view expressed in the Nuclear Fusion paper Fitzpatrick:2026 can be gained from the “private communication” in a September 10, 2025 email: “a tokamak plasma is quiescent. There are no large amplitude tearing modes, and the equilibrium profiles are determined by transport rather than MHD activity. The only exception is the sawtooth oscillation, which periodically generates a short burst of MHD activity that prevents the central $q$ value from falling significantly below unity. Given that $q(0)$ is pinned to a value close to unity, and $q(a)$ is fixed by the plasma current, the average broadness of the current profile is easily controlled. Experimentalists have also become adept at further tailoring the current profile by means of off-axis heating. It is not true that there is an inevitable drift to a disruptive state in a tokamak plasma. In fact, tokamak discharges could be maintained indefinitely were it not for the limit in the flux-swing of the central solenoid, and the gradual build up of impurities.” This reads as if disruptions can be easily avoided by a competent machine operator, but this was not the case in JET shutdowns Sozzi:2020 .
Although the arXiv article Boozer:flux did not imply that tokamaks have “an inevitable drift to a disruptive state,” ensuring tokamaks have a sufficiently low disruption rate for the economic feasibility of power plants is non-trivial and requires improved understanding. Ensuring disruptions are sufficiently rare for the feasibility of tokamak power plants would seem to require the identification of the externally controllable plasma parameters that ensure the plasma is safe against disruptions. This requirement is more stringent in power plants than in existing tokamaks or in JET because of the limitations on diagnostics and actuators. Generative AI using data from both experiments and simulations could help in this identification.
Empirical studies are essential for identifying parameter regions of low disruptivity. Studies based on the DCAF code developed by Steve Sabbagh’s group have been used to determine Sabbagh:2024 the empirical dependence of tokamak disruptions on plasma quantities such as the internal inductance $\ell_{i}$ and edge safety factor $q_{95}$, Figure 2, The results for MAST-U, which has highly non-circular magnetic surfaces, gave preliminary evidence that an internal inductance that is either too low or too high gives disruptive plasmas with a stable region between. An empirical study that is related to that of Figure 2 is given in Figure 3, which is a study of disruptions in JET with an ITER-like wall, JET-dis:2020 . That figure shows that internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ are not sufficient to define a stable region. Are the numerical simulations of tokamaks sufficiently sophisticated and reliable to explain Figures 2 and 3?
Figure 2: MHD-caused disruptions are illustrated in an internal inductance $\ell_{i}$ and edge safety factor $q_{95}$ diagram with separate symbols for KSTAR and NSTX-U as well as for two operational years of MAST-U. Two branches of DCS trigger points can be recognized in both MAST-U experimental years. The two branches (visualized with green lines) are separated along the $\ell_{i}$ axis into two regions, with the internal inductance being roughly restricted to $1.5\lesssim\ell_{i}\lesssim 1.7$ in branch 1 and $0.8\lesssim\ell_{i}\lesssim 1.2$ in branch 2 . The region between the branches is free of MHD-caused disruptions for $q_{95}\gtrsim 6.3.$ This is Figure 7 in the paper V. Zamkovska, S.A. Sabbagh, M. Tobin, et al, Nucl. Fusion 64, 066030 (2024).
Figure 3: Pre-disruptive parameters in JET with an ITER-like wall are shown in a $\ell_{i}-q_{95}$ stability diagram. This was Figure 13 in Gerasimov et al, Nucl. Fusion 60, 066028 (2020). As the authors note: “It may be expected that a disruption free space may be defined in the $\ell_{i}-q_{95}$ empirical stability diagram, assuming that plasma current profiles tend to maintain itself inside the permissible values. In reality, the JET-ILW pre-disruptive plasma equilibrium parameters create a diffused cloud on the $\ell_{i}-q_{95}$ stability diagram without room for non-disruptive plasmas.”
A major determinant of disruptivity is the profile of the net plasma current $I(\psi_{t})$ enclosed by a magnetic surface that also encloses a toroidal magnetic flux $\psi_{t}$. The derivative $I^{\prime}(\psi_{t})\equiv dI/d\psi_{t})$ is given by the divergence-free part of $\vec{j}_{||}=(j_{||}/B)\vec{B}$, which is the net plasma current, and is determined by the spatial constancy of the loop voltage. Despite the view expressed in the September 10, 2025 “private communication” quoted above, the naturally arising profiles of $I(\psi_{t})$ during startup, flattop, and shutdown have many uncertainties. Figures 2 and 3 show that actual tokamak plasmas have many values of $\ell_{i}$, which quantifies the peakedness of the current profile, for a given edge safety factor. There is not just the one current profile that the Nuclear Fusion paper Fitzpatrick:2026 focussed on.
What theory can do is identify controllable features in $I(\psi_{t})$ that determine the internal inductance $\ell_{i}$. When the central safety factor $q_{0}$ is held to unity by sawteeth, the derivative $I^{\prime}(0)$ is defined within a factor determined by the ellipticity of the near-axis magnetic surfaces. The determination of the current density $I^{\prime}(\psi_{t})$ by a spatially-constant loop voltage implies $I^{\prime}(\psi_{t})\geq 0$ without negative external current drive. Both a positive $I^{\prime}$ near the edge, as in a H-mode plasma, and triangular shaping of the magnetic surfaces reduce $\ell_{i}$ for a fixed edge to central $q$-factor ratio with an otherwise fixed $I(\psi_{t})$. It is important to determine if extending the empirical data used to construct Figures 2 and 3 from a two parameter to a four parameter fit would define regions that have a sufficiently robust stability against disruptions.
II Poloidal flux and its evolution in tokamaks
II.1 Basic properties the poloidal flux
Although ignored in the Nuclear Fusion paper Fitzpatrick:2026 , fundamental properties of the poloidal flux and its evolution were illustrated in Figure 1, from the arXiv article Boozer:flux . Just below that figure, the properties of the evolution of the poloidal flux enclosed by the circular magnetic axis of a tokamak were obtained by a simple application of Stokes’ Theorem to Faraday’s Law.
The Faraday’s Law result is
$\frac{\partial}{\partial t}\int\vec{B}\cdot d\vec{a}$ $=$ $-\int\left(\vec{\nabla}\times\vec{E}\right)\cdot d\vec{a}$ (1)
$\frac{\partial\psi^{ax}_{p}}{\partial t}$ $=$ $V_{\ell}^{ax},$ (2)
The magnetic flux going down through the magnetic axis, which is a circle of radius $R_{ax}$ is
$\psi_{p}^{ax}\equiv-\int\vec{B}\cdot d\vec{a}$ (3)
integrated over the area enclosed by the axis. The loop voltage at the axis is
$V_{\ell}^{ax}$ $\equiv$ $\oint\vec{E}\cdot d\vec{\ell}$ (4)
$=$ $2\pi R_{ax}\eta_{ax}j_{ax}$ (5)
with $\eta_{ax}$ the resistivity at the axis and $j_{ax}$ the current density flowing along the magnetic field line that forms the axis.
The choice of sign for the poloidal flux in $(R,\varphi,Z)$ cylindrical coordinates,
$\psi_{p}(R)$ $\equiv$ $-2\pi\int_{0}^{R}B_{z}RdR,$ (6)
with $Z=0$, may seem odd, but it results in $\psi_{p}$ increasing with distance $r$ from the magnetic axis, which gives a positive safety factor. The distance from the axis is $r\equiv R_{ax}-R$, and $2\pi(R_{ax}-r)dr$ gives the area element for defining the poloidal flux that lies within a distance $r$ of the axis.
$\psi_{p}^{f^{\prime}}$ $\equiv$ $\psi_{p}(R)-\psi_{p}(R_{ax})$ (7)
$=$ $2\pi\int_{0}^{r}B_{z}(R_{ax}-r)dr,\hskip 14.45377pt\mbox{where}\hskip 14.45377pt$ (8)
$\simeq$ $2\pi R_{ax}\int_{0}^{r}B_{\theta}dr\hskip 14.45377pt\mbox{for $r<<R_{ax}$;}\hskip 14.45377pt$ (9)
The poloidal flux $\psi_{p}^{f^{\prime}}$ is related to the definition in the paper Fitzpatrick:2026 , which used a cylindrical $(r,\theta,z)$ model for the poloidal magnetic field in a tokamak.
The definition of poloidal flux in Fitzpatrick:2026 differed in two ways from $\psi_{p}^{f^{\prime}}$ of Equation 9. First, a factor of no fundamental importance due to the non-circularity of the magnetic surfaces was included, which for simplicity will be ignored. The effect of magnetic-surface shape on the rotational transform is given by Equation (237) in Reference NF review . Second and what is of fundamental importance, the paper Fitzpatrick:2026 did not include the poloidal flux that lies outside of the plasma but is produced by the plasma current; only the poloidal flux internal to the plasma was retained,
$\psi_{p}^{in}=2\pi R_{ax}\int_{0}^{a}B_{\theta}dr,$ (10)
where $a$ is the plasma radius. An additive time dependent value of the poloidal flux at the magnetic axis, $\psi_{p}^{ax}(t)$, was included in the poloidal flux evolution, see Equation (14), but the time dependence of the poloidal flux at the axis, unlike the poloidal flux in the central solenoid is not directly controlable.
For the circular surfaces of a $(r,\theta,z)$ model, $B_{\theta}=\mu_{0}I(r)/(2\pi r)$, where $I(r)=2\pi\int_{0}^{r}j_{t}(r)rdr$ and $j_{t}$ is the net toroidal current density. The paper Fitzpatrick:2026 set the plasma contribution to the poloidal flux equal to zero for $r>a$. In a cylindrical approximation, the poloidal flux that lies between the radius of the plasma $a$ and an outer radius $b>a$ is $\{\mu_{0}I_{p}/(2\pi)\}\int_{a}^{b}dr/r$, which is proportional to $\ln(b/a)$ and becomes logarithmically infinite for $b>>a$.
As noted in the arXiv article Boozer:flux and stated in many other places, the poloidal flux of a large aspect ratio toroidal shell with major and minor radii $R$ and $a$ carrying a current $I_{p}$ is
$\psi_{p}^{ex}=\mu_{0}R\left\{\ln\left(\frac{8R}{a}\right)-2\right\}I_{p}.$ (11)
This expression can be obtained using the method of the solved problem 5.32 in Reference Jackson .
For standard tokamaks, the external part of the poloidal flux produced by the plasma current is larger than the flux given by the plasma current that lies within the plasma, the part $\psi_{p}^{in}$ of Equation (10). Equation (21) of the Nuclear Fusion paper Fitzpatrick:2026 states the typical poloidal flux is $\psi_{0}=2\pi R_{ax}aB_{\theta}(a)$. When one ignores the increased current required for a given flux due to non-circularity, which affects both $\psi_{p}^{in}$ and $\psi_{p}^{ex}$, Fitzpatrick’s typical flux is $\psi_{0}=\mu_{0}R_{ax}I_{p}$. The ratio $\psi_{p}^{ex}/\psi_{0}=\ln(8R/a)-2\approx 1.178$ when $R/a=3$. For a parabolic current profile, the ratio of the external to the internal flux produced by the plasma current is even larger, $\psi_{p}^{ex}/\psi_{p}^{in}=1.57$.
The poloidal flux enclosed by the magnetic axis, Equation (3), has two important parts: the part produced by the plasma current $\Psi_{p}$ and the poloidal flux in the solenoid, $\psi_{sol}(t)$, which is the only directly controllable function of time;
$\psi_{p}^{ax}$ $=$ $\Psi_{p}+\psi_{sol}(t);$ (12)
$\Psi_{p}$ $=$ $-(\psi_{p}^{in}+\psi_{p}^{ex}).$ (13)
In the approximation of circular surfaces and a large aspect ratio, $R/a>>1$, Equation (10) gives $\psi_{p}^{in}$, and Equation (11) gives $\psi_{p}^{ex}$.
II.2 Disentangling the expressions used in Fitzpatrick:2026
The paper Fitzpatrick:2026 entangles arguments about poloidal flux evolution with issues that are in principle separable, such as heat transport. This entanglement produces confusion—apparently in Fitzpatrick:2026 itself—and a disentanglement is required.
The first step in a disentanglement is determining what is the precise definition of the poloidal flux in Fitzpatrick:2026 and what are the equations it obeys in the cylindrical $(r,\theta,z)$ approximation that is utilized. Equations (6) and (9) in Fitzpatrick:2026 together with Equation (37) applied at $\hat{r}\equiv r/a=0$, where $r=a$ is the plasma edge, imply that the definition of the poloidal flux $\psi_{f}(r,t)$ is
$\psi_{f}(r,t)$ $=$ $\psi_{f}(0,t)+2\pi R_{ax}\int_{0}^{r}B_{\theta}dr;$ (14)
$\frac{d\psi_{f}(0,t)}{dt}$ $=$ $2\pi R_{ax}E_{z}$ (15)
$=$ $V_{\ell}^{ax},$ (16)
the loop voltage at the magnetic axis.
The poloidal magnetic flux is the flux that goes down through the central hole of a toroidal magnetic surface that encloses toroidal flux $\psi_{t}\simeq B_{ax}\pi r^{2}$.
The time derivative of the poloidal flux outside a magnetic surface that encloses toroidal flux $\psi_{t}$ satisfies exactly
$\frac{\partial\psi_{p}(\psi_{t},t)}{\partial t}$ $=$ $V_{\ell}(\psi_{t},t);$ (17)
$V_{\ell}(\psi_{t},t)$ $\equiv$ $\lim_{L\rightarrow\infty}\frac{\int_{-L}^{L}\vec{E}\cdot d\vec{\ell}}{\int_{-L}^{L}\vec{\nabla}\left(\frac{\varphi}{2\pi}\right)\cdot d\vec{\ell}}.$ (18)
This is true during any period that the magnetic surface exists, where $d\vec{\ell}$ is the differential distance along a magnetic field line in that surface. Unlike, the equivalent equation at the magnetic axis, where $\psi_{t}=0$, this is not trivially derived from Faraday’s Law and Stokes’ Theorem, but is given by Equations (27) and (31) in Reference Boozer:RMP .
A fundamental mathematical point that was missed in Fitzpatrick:2026 is that the toroidal magnetic flux is defined by an inside-outwards area integral from the magnetic axis while the poloidal flux is defined by an outward-inwards integral towards the magnetic axis from the $R=0$ axis of the torus.
The necessary condition in Fitzpatrick:2026 for a sufficiently slow ramp up or down to avoid a disruption is given by Equation (42), which is equivalent to
$V_{\ell}^{ax}$ $>$ $\left|2\pi R_{ax}\frac{\partial}{\partial t}\int_{0}^{a}B_{\theta}dr\right|\hskip 14.45377pt\mbox{or}\hskip 14.45377pt$ (19)
$>$ $\left|\frac{d\psi_{p}^{in}}{dt}\right|.$ (20)
This condition can be expressed as a condition on the loop voltage at the plasma edge, $V_{\ell}^{edge}=(d\psi_{p}/dt)_{edge}$, which is at $r=a$. The poloidal flux at the plasma edge is given by
$\psi_{p}^{edge}(t)$ $=$ $\psi_{p}^{ax}-\psi_{p}^{in}\hskip 14.45377pt\mbox{or}\hskip 14.45377pt$ (21)
$=$ $\psi_{sol}(t)+\psi_{p}^{ex}.$ (22)
During a ramp up, $d\psi_{p}^{in}/dt>0$, the necessary condition Equation (20), which is a rewrite of Equation (42) in Fitzpatrick:2026 , for avoiding a disruption is $V_{\ell}^{edge}>0$, which is a necessary condition for any ramp up but not a sufficient condition for avoiding a disruption. During a ramp-down, $d\psi_{p}^{in}/dt<0$, the necessary condition of Fitzpatrick:2026 for avoiding a disruption is $V_{\ell}^{edge}<2V_{\ell}^{ax}$, which is also not a sufficient condition for avoiding a disruption. What is required is that the current density profile not be disruptive as discussed in Subsection II.4.
II.3 Non-disruptivity of the current profile of Fitzpatrick:2026
Although the paper Fitzpatrick:2026 gave an expression for the profile of perpendicular energy diffusivity of some generality, Equation (88), the paper actually determined the current profile that would be obtained for only one case, a radially constant energy-diffusivity profile. The resulting current profile is not disruptive using the internal inductance $\ell_{i}$, Equation (90) in Fitzpatrick:2026 , and the edge safety factor $q_{a}$ based on the 1987 Cheng, Furth, and Boozer study MHD stab in a cylindrical approximation of disruptions in TFTR. In the paper Fitzpatrick:2026 , the only source of heating was Ohmic and the only source of cooling was the thermal transport coefficient, which obeyed a strong assumption—a fixed profile in time that was spatially constant. In reality, fusion power is centrally concentrated and quickly ceases during shutdown. The radiated power from impurities has a complicated dependence on temperature, and impurities modify the resistivity profile through the $Z_{eff}$ dependence of the resistivity $\eta$.
What was not done in Fitzpatrick:2026 is to determine the range of profiles of the energy diffusivity that give non-disruptive current profiles by the Cheng et al criterion—even in the approximation of ignoring the subtleties of heating and radiative cooling. Showing stability for one time-independent thermal-diffusivity profile, which was a spatial constant, is not a reliable guide for determining the danger of disruptions.
In a response to an earlier version of this Comment, Fitzpatrick Fitzpatrick:2026A stated: “The transport of electron energy in tokamak plasma has been the subject of extensive research for the last 50 years. There is overwhelming evidence that such transport is diffusive in nature.” Despite the certainty of this statement, the general view among tokamak experts is that energy diffusivity is not always the correct way to characterize energy transport—even when radiative losses are ignored. The temperature profiles can be given rigidity by microturbulence, Holland:2021 ; Turb-profiles:2025 . The study by N.T. Howard et al of microturbulent transport in ITER Turb-profiles:2025 has sufficient information to calculate the loop voltage and determine its spatial constancy. This would determine the consistency of their assumed current profile $I(\psi_{t})$. Much work remains to establish the consistency of current profiles with both thermal transport and stability against disruptions in full ramp-up, flattop, and ramp-down scenarios even as the nuclear power production turns on and off and as impurities gives complicated power losses.
II.4 Actual condition for disruption avoidance
Not only did the paper Fitzpatrick:2026 ignore the simple statements in the arXiv article Boozer:flux about the poloidal flux and its evolution, the paper also ignored the statement in the abstract Boozer:flux about an important result: “A deviation of the profile of the plasma current over its full stability range produces only a small change, $\sim 16\%$, in the poloidal flux produced by the plasma current. This offers a simple explanation of why disruptions in tokamaks are so common, and why current-profile control though difficult seems to be required, especially during shutdown.”
The paper Fitzpatrick:2026 could not examine the correctness of this statement for two reasons. First by ignoring $\psi_{p}^{ex}$, it ignored most of the poloidal flux produced by the plasma current. Second by not considering the range of profiles of current density that are consistent with stability by the Cheng-Furth-Boozer criterion MHD stab , the paper Fitzpatrick:2026 could not estimate what range of poloidal fluxes produced by the plasma current are stable to disruptions.
Only a small fractional change in the poloidal flux is required to move a tokamak equilibrium that is stable to disruptions into a disruptive state. Consequently, a very careful control of the poloidal flux evolution is required during a plasma shut down—all the poloidal flux produced by the plasma current $\Psi_{p}=-(\psi_{p}^{in}+\psi_{p}^{ex})$ must be removed. The length of the timescale for $\Psi_{p}$ removal is so long that the current profile within the plasma has time to relax to nearly its steady-state limit, which is a spatially constant loop voltage $V_{\ell}$ across the plasma. Ignoring current drive and the bootstrap current, as the paper Fitzpatrick:2026 does, the net toroidal current density on each magnetic surface, which is denoted by its enclosed toroidal flux $\psi_{t}$, is
$j_{t}(\psi_{t})=\frac{V_{\ell}}{2\pi R\eta(Z_{eff},T_{e})},$ (23)
where $2\pi R$ is the average distance required for a field line to make a toroidal circuit, $Z_{eff}$ is the correction to the resistivity produced by impurities, and $T_{e}$ is the electron temperature. Neoclassical effects to the resistivity $\eta$ can be easily included. Both the $T_{e}$ and $Z_{eff}$ profiles generally evolve as the plasma evolves and are not independent of time as assumed in Fitzpatrick:2026 . The difficulty of obtaining disruption-free shutdowns in JET was discussed in Sozzi:2020 . This was thought largely due to impurity issues.
In pulsed tokamaks power plants, the pulse length is set by the length of time the central solenoid can supply the loop voltage required to maintain the current. Typically the solenoid is designed to supply a current-maintenance flux comparable to $\Psi_{p}$ during the fusion-production flattop, which implies the current profile of Equation (23) with $V_{\ell}$ essentially the same on all magnetic surfaces in the plasma.
The actual condition for disruption avoidance is that the current profile given by Equation (23), with $V_{\ell}$ the same on all magnetic surfaces, never crosses into a profile unstable to a disruption. This crossing can be estimated using Cheng et al MHD stab , but a far more reliable estimate could be obtained using three-dimensional magnetohydrodynamics codes coupled with comparisons to experiments.
The paper Fitzpatrick:2026 assumed a poloidal field profile that was essentially independent of time, which implies a spatially constant loop voltage in the plasma region. However, a spatially constant loop voltage does not imply a time independent poloidal field profile when the electron temperature and the $Z_{eff}$ profiles are time dependent.
The primary point in the Physics of Plasmas submission, which was the September 9, 2025 version of Boozer:flux , was the spatial constancy of the loop voltage. A major difference in calculations is that the arXiv article Boozer:flux considered a number of current profiles that ranged over the stable region of the Cheng et al MHD stab paper and found the poloidal flux in the plasma changed by relatively small amount over that range of current profiles.
The Nuclear Fusion paper Fitzpatrick:2026 considered only one current profile, which happened to be stable, but this is not a credible argument for the robust stability against disruptions for all realistic current profiles.